OpenStudy (anonymous):

Represent the given decimal as a fraction. 0.8181.....

OpenStudy (anonymous):

81/99

OpenStudy (anonymous):

or 9/11

OpenStudy (anonymous):

1/9 = .111111111111111 2325/9999 = .232523252325

OpenStudy (anonymous):

still confused.

OpenStudy (anonymous):

0.532532532... = 532/999

OpenStudy (anonymous):

0.123456712345671234567...=1234567/9999999 see the pattern yet

OpenStudy (anonymous):

if its over 9, 99, 999, 9999, 99999.... then the number will repeat infinitely

OpenStudy (anonymous):

like 20/99 = .202020202020..

OpenStudy (anonymous):

5/9 = .555555

OpenStudy (anonymous):

oh! now i see it! Thanks :)